OpenStudy (anonymous):

Find the reciprocal of the number. Multiply to check your answer. 2 3/4 a. 2 3/4 b.11/4 c.4/11 d4/23 A. B. C. D. A. B. C. D.

OpenStudy (anonymous):

Mixed fraction is converted to simple fraction by this procedure: \[a \frac{b}{c} = \frac{ac + b}{c}\]

OpenStudy (anonymous):

Likewise : \[2 \frac{3}{4} = ??\]

OpenStudy (anonymous):

\(a=2\), \(b=3\) and \(c=4\) there.

OpenStudy (anonymous):

Correct.